# Non-convergence of the Navier-Stokes equations toward the Euler equations in weak Besov spaces

### v1

@@ 3.1 Construction of Initial Data
We need to introduce smooth, radial cut-off functions to localize the frequency region. Let $\widehat{\phi}\in\mathcal{C}^{\infty}_{0}(\mathbb{R})$ be an even, real-valued and non-negative function on $\mathbb{R}$ and satisfy
@@
For any $p\in[1,\infty]$, there exists two positive constants $c_{1}$ and $C_{1}$, we have
@@
We assume that $p\in[1,\infty)$ without loss of generality. By easy computations, we deduce that
$\widehat{f_{j}}(\xi)=\frac{1}{2}\left[\widehat{\phi}\left(\xi_{1}+\frac{17}{12}2^{j}\right)+\widehat{\phi}\left(\xi_{1}-\frac{17}{12}2^{j}\right)\right]\prod_{i=2}^{d}\widehat{\phi}\left(\xi_{i}\right),$
@@
[…]
@@
where we have used Remark 3.1. We complete the proof of Lemma 3.1. ∎
@@ 3.2 Estimation of Leading Term
Let $u_{0}$ be given by Lemma 3.1. Assume that $(s,p)$ satisfies (1.1). For $t\in(0,1)$, we have for some positive constants $c_{1},C_{1},C_{2}$ independent of $t,\varepsilon$ and $\delta$
@@
From (3.5), then by Bernstein’s inequality (see Lemma 2.1), we obtain from (3.4) that for large $n$
$2^{ns}\big{\|}\Delta_{n}\big{(}e^{t\varepsilon\Delta}u_{0}-u_{0}\big{)}\big{\|}_{L^{p}}$ $\leq t2^{n}2^{ns}\big{\|}{\Delta}_{n}\left(\varepsilon\Delta u_{0}\right)\big{\|}_{L^{p}}$
$\thickapprox t2^{n}2^{ns}\big{\|}{\Delta}_{n}u_{0}\big{\|}_{L^{p}}$
$\thickapprox t\delta 2^{n}.$
@@ 3.3 Estimation of Error Terms
From now on, we set $u^{\rm E}_{t}=u^{\rm E}(t,u_{0})$ for simplicity. Due to the fact $B^{s}_{p,\infty}\hookrightarrow\rm Lip$, we know that there exists a positive time $T=T(\|u_{0}\|_{B^{s}_{p,\infty}})$ such that
$\|u^{\rm E}_{t}\|_{L^{\infty}_{T}B^{s}_{p,\infty}}\leq C\|u_{0}\|_{B^{s}_{p,\infty}}\leq C\delta.$
By the Mean value theorem, from (E), one has
@@
$\left\|u^{\rm E}_{t}-u_{0}+t\mathcal{P}\left(u_{0}\cdot\nabla u_{0}\right)\right\|_{\dot{B}^{s-2}_{p,\infty}}\leq$ $~{}C\int_{0}^{t}\left\|u_{0}\cdot\nabla u_{0}-u^{\rm E}_{\tau}\cdot\nabla u^{\rm E}_{\tau}\right\|_{B^{s-2}_{p,\infty}}\mathrm{d}\tau$
@@
$\leq$ $~{}C\int_{0}^{t}\left\|u^{\rm E}_{\tau}-u_{0}\right\|_{B^{s}_{p,\infty}}\|u_{0}\|_{B^{s}_{p,\infty}}\mathrm{d}\tau$
@@
From now on, we set $u^{\rm NS}_{t}=u^{\rm NS}(t,u_{0})$ for simplicity. Due to the fact $B^{s}_{p,\infty}\hookrightarrow\rm Lip$, we know that there exists a positive time $T=T(\|u_{0}\|_{B^{s}_{p,\infty}})$ such that
$\|u^{\rm NS}_{t}\|_{L^{\infty}_{T}B^{s}_{p,\infty}}\leq C\|u_{0}\|_{B^{s}_{p,\infty}}\leq C\delta.$
@@
$\leq$ $~{}C\int_{0}^{t}\left\|u^{\rm NS}_{\tau}-e^{\tau\varepsilon\Delta}u_{0}\right\|_{B^{s-1}_{p,\infty}}\|u^{\rm NS}_{\tau}\|_{B^{s}_{p,\infty}}\mathrm{d}\tau$
@@ 3.4 Non-Convergence
In this subsection, we shall prove the non-convergence of the solutions for the Navier-Stokes equations toward the Euler equations in the inviscid limit under the framework of weak Besov spaces $B^{s}_{p,\infty}$.
@@
$u^{\rm NS}_{t}-u^{\rm E}_{t}$ $=\left(e^{t\varepsilon\Delta}u_{0}-u_{0}\right)-\int_{0}^{t}e^{(t-\tau)\varepsilon\Delta}\mathcal{P}\left(e^{\tau\varepsilon\Delta}u_{0}\cdot\nabla e^{\tau\varepsilon\Delta}u_{0}-u_{0}\cdot\nabla u_{0}\right)\mathrm{d}\tau$
$\quad-\int_{0}^{t}\left(e^{\tau\varepsilon\Delta}-\mathrm{Id}\right)\mathcal{P}\left(u_{0}\cdot\nabla u_{0}\right)\mathrm{d}\tau-\left[u^{\rm E}_{t}-u_{0}+t\mathcal{P}\left(u_{0}\cdot\nabla u_{0}\right)\right]$
$\quad+u^{\rm NS}_{t}-e^{t\varepsilon\Delta}u_{0}+\int_{0}^{t}e^{(t-\tau)\varepsilon\Delta}\mathcal{P}\left(e^{\tau\varepsilon\Delta}u_{0}\cdot\nabla e^{\tau\varepsilon\Delta}u_{0}\right)\mathrm{d}\tau$
$=\left(e^{t\varepsilon\Delta}u_{0}-u_{0}\right)+\mathbf{I}_{1}+\mathbf{I}_{2}+\mathbf{I}_{3}+\mathbf{I}_{4},$ (3.10)
@@
$\mathbf{I}_{3}:=-\left[u^{\rm NS}_{t}-u_{0}+t\mathcal{P}\left(u_{0}\cdot\nabla u_{0}\right)\right],$
@@
Using the triangle inequality and Proposition 3.2, we have
@@
Thus, choosing $\delta$ sufficiently small, we deduce that
$\left\|u^{\rm NS}_{\varepsilon}(t,u_{0})-u^{\rm E}(t,u_{0})\right\|_{B^{s}_{p,\infty}}$ $\geq c_{0}\delta^{2}.$
Notice that $t\to 0^{+}$ and $\varepsilon\to 0^{+}$ as $n\to\infty$, we complete the proof of Theorem 1.2. $\square$

### v2

@@ 3.1 Construction of Initial Data
We need to introduce smooth, radial cut-off functions to localize the frequency region. Let $\widehat{\phi}\in\mathcal{C}^{\infty}_{0}(\mathbb{R})$ be an even, real-valued and non-negative function on $\mathbb{R}$ that satisfies
@@
We assume that $\widehat{\phi}\in\mathcal{C}^{\infty}_{0}(\mathbb{R})$ is even and real-valued such that $\phi(x)=(2\pi)^{-1}\mathcal{F}^{-1}(\widehat{\phi}(\xi))$ is a real-valued and continuous function on $\mathbb{R}$. It is easy to check that $\|\phi\|_{L^{\infty}}=\phi(0)>0$. Then there exists some $r>0$ such that $\phi(x)\geq\phi(0)/{2}$ for any $x\in B_{r}(0).$ Moreover, we have
@@
For any $p\in[1,\infty]$, there exists two positive constants $c_{1}$ and $C_{1}$, such that
@@
Remark 3.2.
Let $f(x)\in\{\sin x,\cos x\}$. For any $p\in[1,\infty]$, there exists some $r>0$ and positive constant $M$, such that for $n\gg 1$
$\left\|\phi(x)f\left(\frac{17}{12}2^{n}x\right)\right\|_{L^{p}(\mathbb{R})}\geq\frac{\phi(0)}{2}\left\|f\left(\frac{17}{12}2^{n}x\right)\right\|_{L^{p}(B_{r}(0))}\geq M.$
In fact, for the case when $p=\infty$, we have for some $n$ large enough
$\left\|f\left(\frac{17}{12}2^{n}x\right)\right\|_{L^{\infty}(B_{r}(0))}=\left\|f(x)\right\|_{L^{\infty}(B_{\lambda_{n}}(0))}=1\quad\text{with}\quad\lambda_{n}:=\frac{17}{12}r2^{n}.$
For the case when $p\in[1,\infty)$, there exists a positive integer number $N$ such that for $n>N$
$\left\|f\left(\frac{17}{12}2^{n}x\right)\right\|^{p}_{L^{p}(B_{r}(0))}$ $=\frac{2r}{\lambda_{n}}\int^{\lambda_{n}}_{0}|f(x)|^{p}\mathrm{d}x\geq\frac{r}{\pi}\int^{\pi}_{0}|f(x)|^{p}\mathrm{d}x,$
since
$\lim_{n\rightarrow\infty}\frac{1}{\lambda_{n}}\int_{0}^{\lambda_{n}}|f(x)|^{p}\mathrm{d}x=\frac{1}{\pi}\int^{\pi}_{0}|f(x)|^{p}\mathrm{d}x.$
@@
By easy computations, we deduce that
$\widehat{f_{j}}(\xi)=\frac{1}{2}\left[\widehat{\phi}\left(\xi_{1}+\frac{17}{12}2^{j}\right)+\widehat{\phi}\left(\xi_{1}-\frac{17}{12}2^{j}\right)\right]\prod_{i=2}^{d}\widehat{\phi}\left(\xi_{i}\right).$
Due to the support condition of $\widehat{\phi}$, we see that
$\mathrm{supp}\ \widehat{f_{j}}(\xi)\subset\mathcal{S}_{1}\cup\mathcal{S}_{2},$
where
$\mathcal{S}_{1}:=\left\{\xi\in\mathbb{R}^{d}:\left|\xi_{1}+\frac{17}{12}2^{j}\right|\leq\frac{1}{2^{d}},\ |\xi_{i}|\leq\frac{1}{2^{d}},\ 2\leq i\leq d\right\},$
$\mathcal{S}_{2}:=\left\{\xi\in\mathbb{R}^{d}:\left|\xi_{1}-\frac{17}{12}2^{j}\right|\leq\frac{1}{2^{d}},\ |\xi_{i}|\leq\frac{1}{2^{d}},\ 2\leq i\leq d\right\},$
@@
$\Delta_{k}u_{0}:=\delta 2^{-k(s+1)}\left(-\partial_{2}f_{k},\;\partial_{1}f_{k},\;0,\;\cdots,\;0\right),$
@@
where we have used Remark 3.1 and Remark 3.2. We complete the proof of Lemma 3.1. ∎
@@ 3.2 Estimation of Leading Term
Let $u_{0}$ be given by Lemma 3.1. Assume that $(s,p)$ satisfies (1.1). For small $t>0$, we have for some positive constants $c_{1},C_{1},C_{2}$ independent of $t,\varepsilon$ and $\delta$
@@
From (3.6) and using the classical $L^{p}$-$L^{p}$ estimate: $\|e^{\varepsilon\tau\Delta}f\|_{L^{p}}\leq\|f\|_{L^{p}}$ with $1\leq p\leq\infty$, we obtain that for large $n$
$2^{ns}\big{\|}\Delta_{n}\big{(}e^{t\varepsilon\Delta}u_{0}-u_{0}\big{)}\big{\|}_{L^{p}}$ $\leq t2^{ns}\big{\|}{\Delta}_{n}\left(\varepsilon\Delta u_{0}\right)\big{\|}_{L^{p}}\leq Ct\delta 2^{n}.$
In fact, for large $n$, by Bernstein’s inequality (see Lemma 2.1) and from (3.5), we have
$2^{ns}\big{\|}{\Delta}_{n}\left(\varepsilon\Delta u_{0}\right)\big{\|}_{L^{p}}\thickapprox 2^{-n}2^{2n}2^{ns}\big{\|}{\Delta}_{n}u_{0}\big{\|}_{L^{p}}\thickapprox\delta 2^{n}.$
@@ 3.3 Estimation of Error Terms
Let $u_{0}$ be given by Lemma 3.1. Assume that $(s,p)$ satisfies (1.1). For small $t>0$, we have for some positive constant $C$ independent of $t,\varepsilon$ and $\delta$
@@
From now on, we set $u^{\rm E}_{t}=u^{\rm E}(t,u_{0})$ for simplicity. By [1, Theorem 7.1], we know that there exists a positive time $T_{1}=T_{1}(\|u_{0}\|_{B^{s}_{p,\infty}})$ such that (E) has a unique solution $u^{\rm E}_{t}\in L^{\infty}_{T_{1}}B^{s}_{p,\infty}$ which satisfies
$\|u^{\rm E}_{t}\|_{L^{\infty}_{T_{1}}B^{s}_{p,\infty}}\leq C\|u_{0}\|_{B^{s}_{p,\infty}}\leq C\delta.$
By the Fundamental Theorem of Calculus, from (E), one has
@@
$\left\|u^{\rm E}_{t}-u_{0}+t\mathcal{P}\left(u_{0}\cdot\nabla u_{0}\right)\right\|_{\dot{B}^{s-2}_{p,\infty}}\leq$ $~C\int_{0}^{t}\left\|u_{0}\cdot\nabla u_{0}-u^{\rm E}_{\tau}\cdot\nabla u^{\rm E}_{\tau}\right\|_{\dot{B}^{s-2}_{p,\infty}}\mathrm{d}\tau$
@@
$\leq$ $~C\int_{0}^{t}\left\|u^{\rm E}_{\tau}-u_{0}\right\|_{B^{s-1}_{p,\infty}}\left(\|u_{0}\|_{B^{s}_{p,\infty}}+\|u^{\rm E}_{\tau}\|_{B^{s}_{p,\infty}}\right)\mathrm{d}\tau$
@@
Let $u_{0}$ be given by Lemma 3.1. Assume that $(s,p)$ satisfies (1.1). For small $t>0$, we have for some positive constant $C$ independent of $t,\varepsilon$ and $\delta$
@@
From now on, we set $u^{\rm NS}_{t}=u^{\rm NS}_{\varepsilon}(t,u_{0})$ for simplicity. By [16, Remark 2], we know that there exists a positive time $T_{2}\geq T_{1}$ such that
$\|u^{\rm NS}_{t}\|_{L^{\infty}_{T_{2}}B^{s}_{p,\infty}}\leq C\|u_{0}\|_{B^{s}_{p,\infty}}\leq C\delta.$
@@
$\leq$ $~C\int_{0}^{t}\left\|u^{\rm NS}_{\tau}-e^{\tau\varepsilon\Delta}u_{0}\right\|_{B^{s-1}_{p,\infty}}\left(\|u_{0}\|_{B^{s}_{p,\infty}}+\|u^{\rm NS}_{\tau}\|_{B^{s}_{p,\infty}}\right)\mathrm{d}\tau$
@@ 3.4 Non-Convergence
In this subsection, we shall prove the non-convergence of the Navier-Stokes equations toward the Euler equations in the inviscid limit under the framework of endpoint Besov spaces $B^{s}_{p,\infty}$.
@@
$u^{\rm NS}_{t}-u^{\rm E}_{t}$ $=\left(e^{t\varepsilon\Delta}u_{0}-u_{0}\right)+\mathcal{E},$ (3.11)
where
$\mathcal{E}:=-\int_{0}^{t}e^{(t-\tau)\varepsilon\Delta}\mathcal{P}\left(u^{\rm NS}_{\tau}\cdot\nabla u^{\rm NS}_{\tau}\right)\mathrm{d}\tau+\int_{0}^{t}\mathcal{P}\left(u^{\rm E}_{\tau}\cdot\nabla u^{\rm E}_{\tau}\right)\mathrm{d}\tau.$
We decompose the error term $\mathcal{E}$ into four terms
$\mathcal{E}$ $=-\int_{0}^{t}e^{(t-\tau)\varepsilon\Delta}\mathcal{P}\left(e^{\tau\varepsilon\Delta}u_{0}\cdot\nabla e^{\tau\varepsilon\Delta}u_{0}-u_{0}\cdot\nabla u_{0}\right)\mathrm{d}\tau$
$\quad-\int_{0}^{t}\left(e^{\tau\varepsilon\Delta}-\mathrm{Id}\right)\mathcal{P}\left(u_{0}\cdot\nabla u_{0}\right)\mathrm{d}\tau+\int_{0}^{t}\mathcal{P}\left(u^{\rm E}_{\tau}\cdot\nabla u^{\rm E}_{\tau}-u_{0}\cdot\nabla u_{0}\right)\mathrm{d}\tau$
$\quad+\int_{0}^{t}e^{(t-\tau)\varepsilon\Delta}\mathcal{P}\left(e^{\tau\varepsilon\Delta}u_{0}\cdot\nabla e^{\tau\varepsilon\Delta}u_{0}-u^{\rm NS}_{\tau}\cdot\nabla u^{\rm NS}_{\tau}\right)\mathrm{d}\tau,$
or equivalently,
$\mathcal{E}=\sum_{i=1}^{4}\mathbf{I}_{i},$ (3.12)
@@
$\mathbf{I}_{3}:=-\left[u^{\rm E}_{t}-u_{0}+t\mathcal{P}\left(u_{0}\cdot\nabla u_{0}\right)\right],$
@@
From (3.11) and (3.12), using the triangle inequality and Proposition 3.1, we have
@@
Then, choosing $\delta$ sufficiently small, we deduce that
$\left\|u^{\rm NS}_{\varepsilon}(t,u_{0})-u^{\rm E}(t,u_{0})\right\|_{B^{s}_{p,\infty}}$ $\geq c_{0}\delta^{2}=:\eta_{0}>0.$
Thus we complete the proof of (1.2) in Theorem 1.2. Notice that $t\to 0^{+}$ and $\varepsilon\to 0^{+}$ as $n\to\infty$, one has
$\liminf_{\varepsilon\downarrow 0}\left\|u^{\rm NS}_{\varepsilon}(t,u_{0})-u^{\rm E}(t,u_{0})\right\|_{B^{s}_{p,\infty}}\geq\eta_{0}>0,$
which implies the non-convergence of Theorem 1.2. In fact, if $\lim\limits_{\varepsilon\downarrow 0}\left\|u^{\rm NS}_{\varepsilon}(t,u_{0})-u^{\rm E}(t,u_{0})\right\|_{L_{T}^{\infty}B^{s}_{p,\infty}}=0$ were true for any small $T$, this will lead to a contradiction. $\square$
Acknowledgments
The authors would like to express their gratitude to the anonymous referees for valuable suggestions and comments which greatly improved the paper.